# Collision-Free Navigation of Wheeled Mobile Robots: An Integrated Path Planning and Tube-Following Control Approach

### v1

@@ VI Conclusion
$\dot{W}_{i}(\mathbf{x})=-k_{0}(\mathbf{x})(1-\phi(\text{d}_{\mathcal{O}}(\mathbf{x})))(\mathbf{x}-\mathbf{s}_{i})^{\top}(\mathbf{x}-\mathbf{x}^{\ast}).$ (60)
Since $(\mathbf{x}-\mathbf{s}_{i})^{\top}(\mathbf{x}-\mathbf{x}^{\ast})\geq 0$ for all $\mathbf{x}\in\mathcal{M}_{i}$, one can conclude from (60) that $\dot{W}_{i}(\mathbf{x})\leq 0$ on the set $\mathcal{M}_{i}$ and $\dot{W}_{i}(\mathbf{x})=0$ only occurs at $\mathbf{x}=\mathbf{s}_{i}$. This implies that $\mathcal{M}_{i}$ is a stable manifold of $\mathbf{s}_{i}$ and any initial state on it will converge to $\mathbf{s}_{i}$. Thus, all $\mathbf{x}(0)\in\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})\cap\mathcal{M}_{i}\subset\mathcal{M}_{i}$ will converge to $\mathbf{s}_{i}$. Case 2): $\mathbf{x}(0)\in\mathcal{N}_{i}\setminus\{\mathbf{s}_{i}\}$. For all $\mathbf{x}\in\mathcal{N}_{i}\setminus\{\mathbf{s}_{i}\}$, we have $\text{d}_{\mathcal{O}}(\mathbf{x})=\epsilon$ and $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))=1$. As such, (55) becomes
$\dot{\mathbf{x}}=-k_{0}(\mathbf{x})\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast}),$ (61)
@@
$\dot{W}_{i}(\mathbf{x})=$ $~{}k_{0}(\mathbf{x})\alpha_{i}(\mathbf{x}-\mathbf{x}^{\ast})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast})$
$-k_{0}(\mathbf{x})(1-\alpha_{i})(\mathbf{x}-\mathbf{c}_{i})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast})$
$=$ $~{}k_{0}(\mathbf{x})\alpha_{i}(\mathbf{x}-\mathbf{x}^{\ast})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast})$
$=$ $~{}k_{0}(\mathbf{x})\alpha_{i}(1-\cos\psi)\|\mathbf{x}-\mathbf{x}^{\ast}\|^{2}>0,$ (62)
@@
$\dot{V}_{i}(\mathbf{x})=-k_{0}(\mathbf{x})(\mathbf{x}-\mathbf{c}_{i})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast})=0,$ (64)
showing that for all $\mathbf{x}\in\mathcal{N}_{i}\setminus\{\mathbf{s}_{i}\}$, the control vector is tangent to the boundary of $\mathcal{O}_{i}^{\epsilon}$, thus directing the robot to move along $\partial\mathcal{O}_{i}^{\epsilon}$. In view of (VI) and (64), we can conclude that if $\mathbf{x}(0)\in\mathcal{N}_{i}\setminus{\mathbf{s}_{i}}$, then $\mathbf{x}$ will keep away from $\mathbf{s}_{i}$ along $\partial\mathcal{O}_{i}^{\epsilon}$ until it leaves the set $\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$. Case 3): $\mathbf{x}(0)\in\mathcal{F}_{i}\setminus(\mathcal{M}_{i}\cup\mathcal{N}_{i})$. Actually, for any $\mathbf{x}\in\mathcal{F}_{i}\setminus(\mathcal{M}_{i}\cup\mathcal{N}_{i})$, it holds that $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))\in(0,1)$, and $\bf{\kappa}_{0}(\mathbf{x})$ and $\mathbf{b}(\mathbf{x})$ are not collinear. Moreover, we can decompose $\bm{\tau}(\mathbf{x})=\mathbf{\Pi}(\mathbf{x})\bm{\kappa}{0}(\mathbf{x})$ into two components: the component $\bm{\tau}(\mathbf{x})_{\parallel}=\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x})\bm{\tau}(\mathbf{x})$ parallel to $\mathbf{b}(\mathbf{x})$ and the component $\bm{\tau}(\mathbf{x})_{\perp}=(\mathbf{I}_{2}-\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x}))\bm{\tau}(\mathbf{x})$ perpendicular to $\mathbf{b}(\mathbf{x})$, as shown in Fig. 13. With (19) in mind, one can verify that
@@
which implies that $\bm{\tau}(\mathbf{x}){\parallel}$ and the bearing vector $\mathbf{b}(\mathbf{x})$ have the same direction, causing $\bm{\tau}(\mathbf{x}){\parallel}$ to point towards $\mathbf{c}_{i}$. A straightforward calculation yields $(\mathbf{I}_{2}-\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x})\mathbf{\Pi}(\mathbf{x})=\mathbf{Q}(\mathbf{b}(\mathbf{x}))$, which allows us to express $\bm{\tau}(\mathbf{x})_{\perp}=\mathbf{Q}(\mathbf{b}(\mathbf{x}))\bm{\kappa}_{0}(\mathbf{x})$. Under $\bm{\tau}(\mathbf{x})_{\perp}$, (55) turns to be (61), which results in $\dot{W}_{i}(\mathbf{x})>0$ as seen in (VI). Thus, $\bm{\tau}(\mathbf{x})_{\perp}$ generates a velocity component tangent to $\mathbf{b}$, causing $\mathbf{x}$ to move away from $\mathbf{s}_{i}$. Since $\bm{\tau}(\mathbf{x})$ and the vector $\mathbf{s}_{i}-\mathbf{x}$ are situated on opposite sides of $\mathbf{b}(\mathbf{x})$ (see Fig. 13), $\mathbf{x}$ cannot move toward $\mathbf{s}_{i}$ on the set $\mathcal{F}_{i}\setminus(\mathcal{M}_{i}\cup\mathcal{N}_{i})$. From Fig. 13, one can intuitively observe that $\bm{\tau}(\mathbf{x})$ guides $\mathbf{x}$ towards the boundary of $\mathcal{O}_{i}^{\epsilon}$ (attributed to $\bm{\tau}(\mathbf{x})_{\parallel})$, while simultaneously keeping away from the manifold $\mathcal{M}_{i}$ (attributed to $\bm{\tau}(\mathbf{x})_{\perp}$). Therefore, for any $\mathbf{x}(0)\in\mathcal{F}_{i}\setminus(\mathcal{M}_{i}\cup\mathcal{N}_{i})$, the solution of (11) will either directly leaves the ball $\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$, or first converges to the set $\mathcal{N}_{i}\setminus\{\mathbf{s}_{i}\}$ and then leaves the ball $\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$ along $\partial\mathcal{O}_{i}^{\epsilon}$ (as shown in Case 2). The three cases above demonstrate that each stationary point $\mathbf{s}_{i}$ ($i\in\mathbb{I}$) is an unstable fixed point, but there exists one line of initial conditions, namely the stable manifold $\mathcal{M}_{i}$, that is attracted to $\mathbf{s}_{i}$. Note that $\mathcal{M}_{i}$ is a 1-dimensional manifold with boundary and thus has zero measure [34]. As such, $\mathbf{x}^{\ast}$ is an almost globally asymptotically stable equilibrium, with a basin of attraction consisting of the free space $\mathcal{X}_{\epsilon}$, except for a set of measure zero. Furthermore, as $\mathbf{x}^{\ast}\in\text{int}(\mathcal{X}_{\epsilon})$, there certainly exists $r^{\ast}>0$ such that $\text{d}_{\mathcal{O}}(\mathbf{x})>\epsilon^{\ast}$ for all $\mathbf{x}\in\mathcal{B}(\mathbf{x}^{\ast},r^{\ast})$. On the ball $\mathcal{B}(\mathbf{x}^{\ast},r^{\ast})$, the kinematics reduces to
$\dot{\mathbf{x}}=-k_{0}(\mathbf{x})(\mathbf{x}-\mathbf{x}^{\ast}),$ (66)
It can be easily verified that $k_{0}(\mathbf{x})\geq\alpha/\sqrt{(r^{\ast})^{2}+\beta^{2}}$ on the ball $\mathcal{B}(\mathbf{x}^{\ast},r^{\ast})$. Given this, (66) implies the local exponential stability of $\mathbf{x}=\mathbf{x}^{\ast}$. Furthermore, by using the inequality $|z|-z^{2}/\sqrt{z^{2}+\beta^{2}}\leq\beta$, $\forall z\in\mathbb{R}$, (52) reduces to
$\dot{W}_{i}(\mathbf{x})\leq$ $-\alpha\|\mathbf{x}-\mathbf{x}^{\ast}\|+\alpha\beta$
$\leq$ $-\sqrt{2}\alpha W_{i}(\mathbf{x})^{\frac{1}{2}}+\alpha\beta,$ (67)

### v2

@@ VI Conclusion
$\dot{W}_{i}(\mathbf{x})=-k_{0}(1-\phi(\text{d}_{\mathcal{O}}(\mathbf{x})))(\mathbf{x}-\mathbf{s}_{i})^{\top}(\mathbf{x}-\mathbf{x}^{\ast}).$ (52)
Since $(\mathbf{x}-\mathbf{s}_{i})^{\top}(\mathbf{x}-\mathbf{x}^{\ast})\geq 0$ for all $\mathbf{x}\in\mathcal{M}_{i}$, one can conclude from (52) that $\dot{W}_{i}(\mathbf{x})\leq 0$ on the set $\mathcal{M}_{i}$ and $\dot{W}_{i}(\mathbf{x})=0$ only occurs at $\mathbf{x}=\mathbf{s}_{i}$. This implies that $\mathcal{M}_{i}$ is a stable manifold of $\mathbf{s}_{i}$ and any initial state on it will converge to $\mathbf{s}_{i}$. Thus, all $\mathbf{x}(0)\in\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})\cap\mathcal{M}_{i}\subset\mathcal{M}_{i}$ will converge to $\mathbf{s}_{i}$.
Case 2): $\mathbf{x}(0)\in\mathcal{N}_{i}\setminus\{\mathbf{s}_{i}\}$. For all $\mathbf{x}\in\mathcal{N}_{i}\setminus\{\mathbf{s}_{i}\}$, we have $\text{d}_{\mathcal{O}}(\mathbf{x})=\epsilon$ and $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))=1$. As such, (47) becomes
$\dot{\mathbf{x}}=-k_{0}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast}),$ (53)
@@
$\dot{W}_{i}(\mathbf{x})=$ $~{}k_{0}\alpha_{i}(\mathbf{x}-\mathbf{x}^{\ast})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast})$
$-k_{0}(1-\alpha_{i})(\mathbf{x}-\mathbf{c}_{i})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast})$
$=$ $~{}k_{0}\alpha_{i}(\mathbf{x}-\mathbf{x}^{\ast})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast})$
$=$ $~{}k_{0}\alpha_{i}(1-\cos\psi)\|\mathbf{x}-\mathbf{x}^{\ast}\|^{2}>0,$ (54)
@@
$\dot{V}_{i}(\mathbf{x})=-k_{0}(\mathbf{x}-\mathbf{c}_{i})^{\top}\mathbf{Q}(\mathbf{b}(\mathbf{x}))(\mathbf{x}-\mathbf{x}^{\ast})=0,$ (56)
showing that for all $\mathbf{x}\in\mathcal{N}_{i}\setminus\{\mathbf{s}_{i}\}$, the control vector is tangent to the boundary of $\mathcal{O}_{i}^{\epsilon}$, thus directing the robot to move along $\partial\mathcal{O}_{i}^{\epsilon}$. In view of (VI) and (56), we can conclude that if $\mathbf{x}(0)\in\mathcal{N}_{i}\setminus{\mathbf{s}_{i}}$, then $\mathbf{x}$ will keep away from $\mathbf{s}_{i}$ along $\partial\mathcal{O}_{i}^{\epsilon}$ until it leaves the set $\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$.
Case 3): $\mathbf{x}(0)\in\mathcal{F}_{i}\setminus(\mathcal{M}_{i}\cup\mathcal{N}_{i})$. Actually, for any $\mathbf{x}\in\mathcal{F}_{i}\setminus(\mathcal{M}_{i}\cup\mathcal{N}_{i})$, it holds that $\phi(\text{d}_{\mathcal{O}}(\mathbf{x}))\in(0,1)$, and $\bf{\kappa}_{0}(\mathbf{x})$ and $\mathbf{b}(\mathbf{x})$ are not collinear. Moreover, we can decompose $\mathbf{h}(\mathbf{x})=\mathbf{\Pi}(\mathbf{x})\bm{\kappa}_{0}(\mathbf{x})$ into two components: the component $\mathbf{h}(\mathbf{x})_{\parallel}=\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x})\mathbf{h}(\mathbf{x})$ parallel to $\mathbf{b}(\mathbf{x})$ and the component $\mathbf{h}(\mathbf{x})_{\perp}=(\mathbf{I}_{2}-\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x}))\mathbf{h}(\mathbf{x})$ perpendicular to $\mathbf{b}(\mathbf{x})$, as shown in Fig. 13. With (12) in mind, one can verify that
@@
which implies that $\mathbf{h}(\mathbf{x})_{\parallel}$ and the bearing vector $\mathbf{b}(\mathbf{x})$ have the same direction, causing $\mathbf{h}(\mathbf{x})_{\parallel}$ to point towards $\mathbf{c}_{i}$. A straightforward calculation yields $(\mathbf{I}_{2}-\mathbf{b}(\mathbf{x})\mathbf{b}^{\top}(\mathbf{x})\mathbf{\Pi}(\mathbf{x})=\mathbf{Q}(\mathbf{b}(\mathbf{x}))$, which allows us to express $\mathbf{h}(\mathbf{x})_{\perp}=\mathbf{Q}(\mathbf{b}(\mathbf{x}))\bm{\kappa}_{0}(\mathbf{x})$. Under $\mathbf{h}(\mathbf{x})_{\perp}$, (47) turns to be (53), which results in $\dot{W}_{i}(\mathbf{x})>0$ as seen in (VI). Thus, $\mathbf{h}(\mathbf{x})_{\perp}$ generates a velocity component tangent to $\mathbf{b}$, causing $\mathbf{x}$ to move away from $\mathbf{s}_{i}$. Since $\mathbf{h}(\mathbf{x})$ and the vector $\mathbf{s}_{i}-\mathbf{x}$ are situated on opposite sides of $\mathbf{b}(\mathbf{x})$ (see Fig. 13), $\mathbf{x}$ cannot move toward $\mathbf{s}_{i}$ on the set $\mathcal{F}_{i}\setminus(\mathcal{M}_{i}\cup\mathcal{N}_{i})$. From Fig. 13, one can intuitively observe that $\mathbf{h}(\mathbf{x})$ guides $\mathbf{x}$ towards the boundary of $\mathcal{O}_{i}^{\epsilon}$ (attributed to $\mathbf{h}(\mathbf{x})_{\parallel})$, while simultaneously keeping away from the manifold $\mathcal{M}_{i}$ (attributed to $\mathbf{h}(\mathbf{x})_{\perp}$). Therefore, for any $\mathbf{x}(0)\in\mathcal{F}_{i}\setminus(\mathcal{M}_{i}\cup\mathcal{N}_{i})$, the solution of (7) will either directly leaves the ball $\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$, or first converges to the set $\mathcal{N}_{i}\setminus\{\mathbf{s}_{i}\}$ and then leaves the ball $\mathcal{B}(\mathbf{s}_{i},\epsilon^{\prime})$ along $\partial\mathcal{O}_{i}^{\epsilon}$ (as shown in Case 2).
The three cases above demonstrate that each stationary point $\mathbf{s}_{i}$ ($i\in\mathbb{I}$) is an unstable fixed point, but there exists one line of initial conditions, namely the stable manifold $\mathcal{M}_{i}$, that is attracted to $\mathbf{s}_{i}$. Note that $\mathcal{M}_{i}$ is a 1-dimensional manifold with boundary and thus has zero measure [36]. As such, $\mathbf{x}^{\ast}$ is an almost globally asymptotically stable equilibrium, with a basin of attraction consisting of the free space $\mathcal{X}_{\epsilon}$, except for a set of measure zero. Furthermore, as $\mathbf{x}^{\ast}\in\text{int}(\mathcal{X}_{\epsilon})$, there certainly exists $r^{\ast}>0$ such that $\text{d}_{\mathcal{O}}(\mathbf{x})>\epsilon^{\ast}$ for all $\mathbf{x}\in\mathcal{B}(\mathbf{x}^{\ast},r^{\ast})$. On the ball $\mathcal{B}(\mathbf{x}^{\ast},r^{\ast})$, the kinematics reduces to
$\dot{\mathbf{x}}=-k_{0}(\mathbf{x}-\mathbf{x}^{\ast}),$ (58)